OpenStudy (anonymous):

let f(x)=2x-x^2... what is the instantaneous rate of change of f at x=3

OpenStudy (anonymous):

take the derivative of f(x) at x = 3 like this f'(x) = 2 - 2x so f'(3) = 2 - 2(3) = -4

OpenStudy (anonymous):

The istantaneous rate of change at x = 3 is just the value of the derivative at x = 3.

OpenStudy (anonymous):

thank you dudes y'all are the best